OpenStudy (abbycross167):

Will someone please help me? I really need it Graph the following equations using the point-slope method. Graph the point from the equation in yellow, your other points in red and your line in blue. y - 7 = -3(x + 2)

OpenStudy (ram231):

well first you have to distribute the -3(x+2) to get rid of the parenthesis. so it would be -3x+-6. so the equation now is y-7=-3x+-6 can you figure out the rest of the equation?

OpenStudy (abbycross167):

umm no ma'am... I'm not very good at math

OpenStudy (ram231):

hmm. so. i can't give you the straight up answer. wish I could. but I will lead you through it okay?

OpenStudy (abbycross167):

yes ma'am thank you

OpenStudy (mathmale):

Will someone please help me? I really need it Graph the following equations using the point-slope method. Graph the point from the equation in yellow, your other points in red and your line in blue. y - 7 = -3(x + 2) Please note that this equation is already in "point-slope" form. See that "x+2?" This means that the x-coordinate of one of the points on this line is -2. See that "y-7?" This means that the y-coordinate of that point on the line is 7. Graph the point (-2,7), please. You can read the slope (m) directly from the given equation. What is it?

OpenStudy (ram231):

So since -3x+-6 is basically like -3x-6 right? well y-7=-3x-6 you need to get the variables on one side of the = sign, while the others are on the other side. so can you do that?

OpenStudy (abbycross167):

y-7-3x=-6?

OpenStudy (abbycross167):

@mathmale does m = -3?

OpenStudy (ram231):

so in y-7=-3x-6, its like adding 7 to the 6. so its off of that side. so now the equation is y=-3x+1 now how do you get the -3x on the other side of the = sign?

OpenStudy (abbycross167):

1-3=2?

OpenStudy (abbycross167):

or would it be -3-1=-2?

OpenStudy (ram231):

Well I basically gave you the equation to use. y=-3x+1... so 1 would be the y intercept, and -3 would be m... in the equation mx+b=c2 so... how I am writing the equation in my explanation is c2=mx+b. the b variable would be the y intercept right?

OpenStudy (abbycross167):

yes ma'am so that would be 1?

OpenStudy (ram231):

yes. and m would be -3

OpenStudy (abbycross167):

ok so on a graph x=-3 and y=1?

OpenStudy (ram231):

yes.

OpenStudy (abbycross167):

ok you're awsome thank you soo very much!! Could you please help me with one more I have to turn it in by 12:00am

OpenStudy (ram231):

sure thing. you can either make a new question or just continue on this.